# MS-ISSM: Objective Quality Assessment of Point Clouds Using Multi-scale Implicit Structural Similarity

### v1

@@ V-C Performance Comparison
WPC-GPCC PLCC 0.841 0.864
SROCC 0.847 0.866
KROCC 0.655 0.731
RMSE 0.193 0.153
WPC-VPCC PLCC 0.482 0.923
SROCC 0.473 0.921
KROCC 0.343 0.771
RMSE 0.190 0.171
MPCCD-GPCC PLCC 0.734 0.935
SROCC 0.673 0.924
KROCC 0.544 0.842
RMSE 0.280 0.216
MPCCD-VPCC PLCC 0.489 0.838
SROCC 0.460 0.801
KROCC 0.288 0.632
RMSE 0.245 0.190
ALL PLCC 0.611 0.886
SROCC 0.566 0.875
KROCC 0.406 0.712
RMSE 0.259 0.172
We compare the average running times across four datasets on an Intel Core i7-8809G CPU @3.10GHz. As shown in Fig. 6, our method achieves superior efficiency, surpassed only by FRSVR. This speed advantage stems from the proposed implicit feature representation, which eliminates the computationally expensive point-to-point matching process. Furthermore, the pre-computed features of the reference point cloud can be reused across different distortion types, avoiding redundant computation. Conversely, GraphSIM and MS-GraphSIM incur high costs due to keypoint sampling and graph construction, while MS-PSSIM is burdened by high-dimensional multi-scale processing. Although simple single-scale metrics like p2p remain computationally light, our method offers a better trade-off between processing speed and multi-scale performance, making it highly suitable for large-scale PCQA.
Figure 6: Time complexity of different PCQA methods on ALL datasets.
The synthesis of all test results indicates that the proposed MS-ISSM successfully aligns the distorted point cloud with the original using implicit structure by corresponding to the feature function coefficient. This alignment is important as it enhances the correspondence of features within point clouds, facilitating a more precise measurement of distortion. By improving feature correspondence, the MS-ISSM aligns and preserves point clouds’ intrinsic geometric and topological attributes, which are critical for accurate distortion measurement.
@@
V-D Ablation Studies
To comprehensively validate the effectiveness of the proposed MS-ISSM framework, we conducted extensive ablation studies. These experiments are designed to investigate the contribution of three key aspects: 1. the multi-modal implicit features, 2. the multi-scale strategy, and 3. the specific architectural components of the ResGrouped-MLP network. The results are analyzed on the combined ALL dataset to ensure statistical reliability.
@@ V-D Ablation Studies
We first investigate the impact of different basis feature functions by evaluating the performance of MS-ISSM using only single feature types: Luma, Chroma, and Curvature. As illustrated in Fig. 7, while single features, particularly Luma, can provide reasonable quality estimations, they exhibit limitations in capturing the full spectrum of perceptual distortions. For instance, geometric distortions are less perceptible in pure color metrics, and vice versa. The fusion of multi-modal features yields superior robustness, confirming that combining geometry and color attributes is essential for aligning with the HVS.
@@
To verify the necessity of the multi-scale hierarchy, we evaluated the performance at individual spatial scales (High, Medium, and Low). As reported in Table IV, although individual scales achieve decent correlations, their performance fluctuates across different datasets due to varying point cloud densities and content characteristics. For example, the High scale performs better on datasets with fine textures, while the Low scale is more effective for global structural distortions. By integrating all three scales, MS-ISSM effectively aggregates local details and global topology, achieving consistent and optimal performance across diverse datasets.
TABLE IV: The Performance comparison of different scales.
Scale Criteria ICIP WPC SJTU M-PCCD High PLCC 0.752 0.837 0.876 0.791 SROCC 0.705 0.842 0.856 0.829 KROCC 0.594 0.651 0.670 0.672 RMSE 0.216 0.376 0.175 0.151 Medium PLCC 0.812 0.827 0.869 0.911 SROCC 0.825 0.834 0.863 0.926 KROCC 0.648 0.637 0.673 0.767 RMSE 0.221 0.349 0.179 0.010 Low PLCC 0.879 0.814 0.788 0.855 SROCC 0.805 0.809 0.749 0.891 KROCC 0.631 0.612 0.555 0.722 RMSE 0.290 0.286 0.198 0.133
@@
Finally, to justify the design rationale of our regression network, we conducted an ablation study by removing or replacing key modules in the ResGrouped-MLP. The comparison results are summarized in Table V. w/o Log-Modulus Transformation: We removed the Log-Modulus preprocessing and used standard Z-score normalization directly on the raw coefficients. As shown in Table V, this resulted in a performance drop, with SROCC decreasing by 0.041. This confirms that the raw statistical features follow a heavy-tailed distribution, and the proposed Log-Modulus transformation effectively suppresses outliers, enabling the network to learn more robust feature representations. w/o Grouped Encoders (Early Concatenation): To validate the ”Split-Transform-Merge” strategy, we replaced the grouped encoders with a standard MLP that concatenates all multi-scale features at the input stage. The results show that the grouped strategy outperforms early concatenation by 0.03 in PLCC. This suggests that processing Luma, Chroma, and Curvature features independently in the early layers prevents information interference, allowing the network to capture distinct distortion patterns for each channel. w/o Attention Mechanism: We removed the Scale-wise Channel Attention blocks. The decline in performance indicates that the attention mechanism plays a crucial role. It allows the model to adaptively recalibrate the importance of different channels (e.g., assigning higher weights to Luma or Low-frequency geometry), thereby better mimicking the varying sensitivities of the HVS. In summary, each component of the MS-ISSM, from implicit feature extraction to the hierarchical regression network, makes a significant contribution to the final prediction accuracy and robustness.
TABLE V: ABLATION STUDY OF THE PROPOSED RESGROUPED-MLP ARCHITECTURE ON THE ”ALL” DATASET.
@@
Model Variant PLCC SROCC KROCC RMSE Full MS-ISSM 0.913 0.914 0.751 0.188 w/o Log-Modulus 0.886 0.873 0.694 0.201 w/o Grouped Encoders 0.883 0.868 0.688 0.204 w/o Attention Block 0.901 0.904 0.717 0.192
  • The testing results of Cross-dataset.
We conducted cross-dataset evaluations to further validate the effectiveness and generalization capability of the proposed method. Given that the ICIP dataset is significantly smaller than the other datasets, it was used as the test set. We trained the model on the SJTU, M-PCCD, and WPC datasets, respectively, and tested the performance of MS-PSSIM on ICIP. As shown by the experimental results in Table VI, models trained on a single dataset generally performed well when validated on the ICIP dataset. These results demonstrate that the proposed method exhibits strong generalization ability and effectiveness.
TABLE VI: The cross-dataset evaluation results (Test Set: ICIP).
Training Set Criteria PLCC SROCC KROCC RMSE WPC 0.818 0.814 0.643 0.195 SJTU 0.880 0.878 0.694 0.159 M-PCCD 0.819 0.838 0.653 0.199

### v2

@@ V-C Performance Comparison
WPC-GPCC PLCC 0.841 0.865
SROCC 0.847 0.864
KROCC 0.655 0.677
RMSE 0.193 0.165
WPC-VPCC PLCC 0.482 0.683
SROCC 0.473 0.638
KROCC 0.343 0.461
RMSE 0.190 0.176
MPCCD-GPCC PLCC 0.734 0.977
SROCC 0.673 0.974
KROCC 0.544 0.866
RMSE 0.280 0.096
MPCCD-VPCC PLCC 0.489 0.921
SROCC 0.460 0.904
KROCC 0.288 0.750
RMSE 0.245 0.117
ALL PLCC 0.611 0.892
SROCC 0.566 0.892
KROCC 0.406 0.718
RMSE 0.259 0.159
V-D Visual Comparisons
To illustrate the perceptual performance of MS-ISSM against existing methods, we present a qualitative comparison in Fig. 6. We selected four representative point clouds subjected to distinct distortion types: biplanebiplane with octree-based compression, sarahsarah with trisoup-based compression, flowerpotflowerpot with Gaussian noise, and bagbag with video-based compression. We then compared the subjective MOS with the predicted scores from the multi-scale feature-based TCDM, the learning-based FRSVR, and our proposed MS-ISSM.
The visual analysis shows that MS-ISSM accurately captures steady quality drops as distortion levels increase. For example, in biplanebiplane with octree-based compression, our method tracks the MOS drop from 0.950 to 0.210 by predicting scores from 0.939 to 0.216, which avoids the consistent underestimation of TCDM and the excessive penalties of FRSVR at medium distortion levels. Furthermore, MS-ISSM is robust to noise, as seen in flowerpotflowerpot with Gaussian noise where TCDM and FRSVR give excessively low scores to small point shifts. In contrast, MS-ISSM uses continuous RBF functions to smooth out these small errors, staying consistent with the HVS. Additionally, in bagbag with video-based compression, FRSVR incorrectly gives a low score of 0.353 to a high-quality model with an MOS of 0.761, causing an incorrect quality ranking. Meanwhile, MS-ISSM correctly tracks the actual quality drop from 0.741 to 0.289. Overall, this visual analysis confirms that MS-ISSM aligns well with human visual perception.
V-E Sensitivity Analysis
We conducted a sensitivity analysis using the color noise (CN) and geometry noise (GN) subsets from the SJTU dataset to evaluate whether the differences in implicit function coefficients well reflect perceptual distortion. The SJTU dataset defines CN as photometric noise applied to the RGB attributes of randomly selected points. Specifically, the noise is injected into 10%, 30%, 40%, 50%, 60%, and 70% of the points, with corresponding intensity limits set to 10, 30, 40, 50, 60, and 70. GN applies a random Gaussian distributed geometric shift to all points. These shifts are bounded within 0.05%, 0.1%, 0.2%, 0.5%, 0.7%, and 1.2% of the bounding box. We computed the difference in the implicit function coefficient (45-dimensional) for the original and distorted point cloud respectively, and evaluated their correlation in PLCC with the perceptual quality measured in subjective MOS. The results are then compared with traditional point-to-point metrics (PSNR-p2p and PSNR-YUV) on the above subsets.
(a) CN
(b) GN
Figure 7: Sensitivity analysis of the RBF coefficient differences.
As shown in Fig. 7, the quantitative results demonstrate that the RBF coefficient differences exhibit a strong correlation with MOS not only at their peak values but across the broader 45-dimensional feature distribution. Under CN, our features achieve a maximum correlation of 0.82, with 11 out of the 30 color-related dimensions explicitly outperforming the traditional PSNR-YUV metric at 0.77, while PSNR-p2p is inherently inapplicable to pure color distortions. Under GN, the coefficients demonstrate even stronger sensitivity with a maximum correlation of 0.95. Furthermore, this superiority is widely distributed across the feature space under GN, where 8 out of the 15 geometry-related dimensions surpass PSNR-p2p at 0.59, and 27 out of the 30 color-related dimensions exceed PSNR-YUV at 0.62. These findings empirically validate that implicit coefficients capture perceptual degradation much more reliably and comprehensively than discrete point matching.
Moreover, to better interpret the superiority of coefficient-space difference, we analyzed soldiersoldier subjected to two different types of distortions: noise and downsampling. As shown in Fig. 8, we compare the predicted normalized quality scores of PSNR-YUV and our MS-ISSM against the subjective MOS. Because the unstructured nature of point clouds leads to local point coordinate shifts in both scenarios, the traditional point-to-point metric (PSNR-YUV) fails to distinguish their actual perceptual differences, producing almost identical quality scores of 0.61 and 0.59 for downsampling and noise, respectively. In contrast, human vision is relatively tolerant to uniform downsampling with a high MOS of 0.77, but highly sensitive to noise-induced degradation with a low MOS of 0.42. Our MS-ISSM effectively captures this perceptual discrepancy, predicting a high quality score of 0.69 for downsampling and appropriately penalizing the noise distortion with a low score of 0.36. Ultimately, the quality predictions of MS-ISSM align consistently with human visual perception.
Figure 8: Coefficient-space difference vs. Point-space error.
V-F Complexity Analysis
We compare the average running times across four datasets on an Intel Core i7-8809G CPU @3.10GHz. As shown in Fig. 9, our method achieves superior efficiency, surpassed only by FRSVR. The efficiency of MS-ISSM comes from replacing global topological comparisons with local matrix operations. When the original point cloud has $N_{\rm{O}}$ points, the sampled reference point set has $N_{\rm{R}}$ points where $N_{\rm{R}}\ll N_{\rm{O}}$, and the local neighborhood size is set to 30. The algorithm processes the data in several steps. First, iterating points to build multi-scale voxel representations takes $\mathcal{O}(N_{\rm{O}})$ time. Next, building a KD-Tree for spatial partitioning requires $\mathcal{O}(N_{\rm{O}}\log N_{\rm{O}})$ operations, and finding nearest neighbors for the $N_{\rm{R}}$ reference points takes $\mathcal{O}(N_{\rm{R}}\log N_{\rm{O}})$ time. For the extraction of curvature features, calculating the local surface variations involves secondary constant-sized neighborhood searches and matrix decompositions within each patch, which securely bounds to an additional ${\mathcal{O}}({N_{\rm{R}}}\log{N_{\rm{O}}})$ time. Subsequently, solving constant-sized linear systems for local RBF fitting adds $\mathcal{O}(N_{\rm{R}})$ time, while processing feature vectors via the ResGrouped-MLP takes $\mathcal{O}(1)$ time. As summarized in Table V, the overall complexity is dominated by space partitioning and neighbor search. Since these operations are performed on both the original and distorted point clouds, the total time doubles but remains bounded by $\mathcal{O}(N_{\rm{O}}\log N_{\rm{O}})$. Graph-based metrics such as GraphSIM and MS-GraphSIM require graph construction or spectral decomposition, while MS-PSSIM and PointPCA are burdened by high-dimensional multi-scale processing in the point space. In contrast, MS-ISSM avoids global topology reconstruction; by restricting complex matrix calculations to highly downsampled reference points, it ensures the linearithmic efficiency shown in Fig. 9. Although simple single-scale metrics like p2p remain computationally light, our method offers a better trade-off between processing speed and multi-scale performance, making it highly suitable for large-scale PCQA.
Figure 9: Time complexity of PCQA methods on four datasets.
TABLE V: complexity analysis of the proposed MS-ISSM.
Processing Stage Complexity multi-scale voxelization $\mathcal{O}(N_{\rm{O}})$ curvature extraction ${\mathcal{O}}({N_{\rm{R}}}\log{N_{\rm{O}}})$ space partitioning $\mathcal{O}(N_{\rm{O}}\log N_{\rm{O}})$ KNN search $\mathcal{O}(N_{\rm{R}}\log N_{\rm{O}})$ RBF implicit fitting $\mathcal{O}(N_{\rm{R}})$ quality regression $\mathcal{O}(1)$ overall complexity $\mathcal{O}(N_{\rm{O}}\log N_{\rm{O}})$
@@ V-F Complexity Analysis
V-G Ablation Studies
To comprehensively validate the effectiveness of the proposed MS-ISSM framework, we conducted extensive ablation studies. These experiments are designed to investigate the contribution of four key aspects: the multi-modal implicit features, the multi-scale strategy, the specific architectural components of the ResGrouped-MLP network, and the key hyperparameters of MS-ISSM. The results are analyzed on the WPC dataset to ensure statistical reliability.
@@ V-G Ablation Studies
We first investigate the impact of different basis feature functions by evaluating the performance of MS-ISSM when removing single feature types: w/o luma, chroma, or curvature. As illustrated in Fig. 10, while the incomplete models can still provide reasonable quality estimations, omitting any specific feature, particularly luma, exhibits limitations in capturing the full spectrum of perceptual distortions. For instance, geometry distortions are less perceptible when structural features are ablated, and vice versa for color. The complete fusion of multi-modal features yields superior robustness, confirming that combining geometry and color attributes is essential for aligning with the HVS.
@@
To verify the necessity of the multi-scale hierarchy, we evaluated the performance by ablating individual spatial scales: w/o High, Medium, or Low. As reported in Table VI, although models missing a single scale still achieve decent correlations, their performance fluctuates across different datasets due to varying point cloud densities and content characteristics. For example, omitting the High scale results in a more significant performance drop on datasets with fine textures, while removing the Low scale more severely impairs the evaluation of global structural distortions. By integrating all three scales, MS-ISSM effectively aggregates local details and global topology, achieving consistent and optimal performance across diverse datasets.
TABLE VI: The Performance comparison of different scales.
Dataset Scale PLCC SROCC KROCC RMSE WPC MS-ISSM (Ours) 0.867 0.859 0.661 11.618 w/o-Scale-H 0.846 0.832 0.642 12.227 w/o-Scale-L 0.833 0.827 0.636 12.674 w/o-Scale-M 0.849 0.839 0.650 12.098
@@
Finally, to justify the design rationale of our regression network, we conducted an ablation study by removing or replacing key modules in the ResGrouped-MLP. The comparison results are summarized in Table VIII. To evaluate the impact of the Log-Modulus transformation, we removed the Log-Modulus preprocessing and used standard Z-score normalization directly on the raw coefficients. As shown in Table VIII, this resulted in a performance drop, with SROCC decreasing by 0.015. This confirms that the raw statistical features follow a heavy-tailed distribution, and the proposed Log-Modulus transformation effectively suppresses outliers, enabling the network to learn more robust feature representations. To validate the ”Split-Transform-Merge” strategy, we replaced the grouped encoders with a standard MLP that concatenates all multi-scale features at the input stage. The results show that the grouped strategy outperforms early concatenation by 0.007 in PLCC. This suggests that processing luma, chroma, and curvature features independently in the early layers prevents information interference, allowing the network to capture distinct distortion patterns for each channel. We also removed the scale-wise channel attention blocks to assess their contribution. The subsequent decline in performance indicates that the attention mechanism plays a crucial role in allowing the model to adaptively recalibrate the importance of different channels, thereby better mimicking the varying sensitivities of the HVS. In summary, each component of the MS-ISSM, from implicit feature extraction to the hierarchical regression network, makes a significant contribution to the final prediction accuracy and robustness.
TABLE VII: ABLATION STUDY OF THE PROPOSED RESGROUPED-MLP ARCHITECTURE ON THE WPC DATASET.
@@
Model Variant PLCC SROCC KROCC RMSE MS-ISSM (Ours) 0.867 0.859 0.661 11.618 w/o Log-Modulus 0.855 0.844 0.657 11.878 w/o Grouped Encoders 0.860 0.848 0.662 11.691 w/o Attention Block 0.852 0.841 0.652 11.987
  • Impact of Key Hyperparameters
We conducted experiments to evaluate the impact of the number of RBF neighbors, multi-scale voxel sizes, the attention bottleneck ratio, and the reference points sampling size. We evaluated these variants on the WPC dataset to ensure reliability, with results summarized in Table VIII. The underlined values denote our chosen configurations.
TABLE VIII: Hyperparameter Ablation on WPC Dataset. The underlined values denote our chosen configurations.
Hyperparameter Configuration PLCC SROCC RBF Neighbors 15 0.839 0.837 30 0.867 0.859 50 0.832 0.825 Voxel Sizes (1.0, 2.0, 4.0) 0.849 0.843 (2.0, 4.0, 8.0) 0.867 0.859 (4.0, 8.0, 16.0) 0.819 0.822 Bottleneck Ratio 2 0.853 0.841 4 0.867 0.859 8 0.846 0.840 Reference Point Size 64 0.833 0.828 32 0.867 0.859 8 0.869 0.860
Regarding the RBF neighbors, we tested values of 15, 30, and 50. A smaller neighborhood of 15 yields a localized receptive field sensitive to isolated noise, while a larger neighborhood of 50 results in over-smoothing that masks fine geometry artifacts and increases computational overhead. Setting the value to 30 balances noise robustness and structural sensitivity. For the multi-scale voxel sizes (H, M, L), we evaluated three different progression ranges: a finer scale of (1.0, 2.0, 4.0), our baseline of (2.0, 4.0, 8.0), and a coarser scale of (4.0, 8.0, 16.0). (1.0, 2.0, 4.0) highlights small details but miss large structural changes. Conversely, (4.0, 8.0, 16.0) downsamples the data too much, losing fine artifacts like color quantization noise.
When analyzing the channel reduction ratio within our attention blocks, we tested values of 2, 4, and 8. A low ratio of 2 retains redundant channel information and increases the risk of overfitting. A high ratio of 8 over-compresses the latent space, causing a loss of semantic interactions between physical attributes including luma, chroma, and curvature.
Finally, we evaluated reference point sampling sizes of 64, 32, and 8. A larger size of 64 leads to insufficient local feature extraction and a noticeable performance drop. A smaller size of 8 yields a marginal performance improvement, but substantially increases the computational burden. Selecting a size of 32 provides the trade-off between computational efficiency and evaluation accuracy. Overall, these empirical validations prove the effectiveness of the architectural choices in MS-ISSM.
V-H Generalization Analysis
To validate the generalization capabilities of our methods, we conducted cross-dataset evaluations using three large-scale benchmark datasets. Specifically, we trained our model entirely on the WPC dataset and tested it directly on the ICIP, SJTU and M-PCCD datasets. To strictly prevent data leakage, the zero-mean normalization mapping fitted on the training set was directly applied to the testing sets. Table LABEL:cross-check below summarizes the results, where SROCC and PLCC values above 0.82 across distinct data distributions can be observed. It is therefore verified that ResGrouped-MLP learns robust and generalized mappings.
TABLE IX: Cross-dataset evaluation results trained on the WPC dataset.
Training Testing PLCC SROCC KROCC RMSE
WPC SJTU 0.840 0.827 0.634 1.298
WPC M-PCCD 0.895 0.918 0.658 0.560
WPC ICIP 0.914 0.918 0.753 0.497